# Rocky planet formation in compact disks around M dwarfs

### v2

@@ 6.1 Outcomes from the formation scenarios
The main conclusion from the different formation scenarios that we explored is the fact that close-in super-Earth formation is possible around M dwarfs in compact disks with low gas disk viscosity () when the core accretion is driven by an efficient pebble radial drift (pebble efficiency between 30 and 60). We found a linear relationship between the super-Earth masses and the stellar mass, in agreement with previous works (e.g. Liu et al., 2019). However, it is not just the isolation mass that sets the final planetary masses, but also the gravitational interactions among planets that lead to collisions among them, thus an increment in their masses. Late collisions among planets become more relevant towards the more massive M dwarfs. As we can see in Section 4, only around a star of 0.5 M⊙ some simulated planets have masses higher than 10 M⊕. Interestingly, such planets had masses below 10 M⊕ while they were immersed in the gas disk. In every case, they reached higher masses after the gas dissipated from the disk, due to late collisions between planets.
@@
For the scenarios in which we study rocky planet formation around a star of 0.1 solar mass, it seems that the formation of Earth-like and super-Earth like planets is independent of the initial location of the planetary embryos, in terms of their final masses and semi-major axis. However, the initial location of the embryos is important in terms of water content, as the scenario in which we considered a sample of embryos initially located inside the snowline is the only one that formed dry planets or planets with less than 10 of water in mass, more consistent with the observations (Luque & Pallé, 2022; Rogers et al., 2023). Even though we did not test the impact of different masses for the initial embryo distribution, we expect that it will not change significantly the final planetary mass range that we found: this is shown by Pan et al. (2024) when studying planet formation around a star of 0.1 M⊙ with , comparing the final planetary masses with both an initial equal mass embryos distribution and an embryo distribution with masses derivate from streaming instability (Liu et al., 2020). However, it would be interesting to see the dependence between the initial mass of the embryos with the final planetary masses in future works.
@@
On the other hand, assuming compact protoplanetary dust disks with higher viscosity () and lower pebble flux is not leading to an efficient planet formation, as all the simulated planets have masses . This is a crucial point as it seems that the migration speed of the embryos immerse in the gas disk as well as the pebble reservoir is directly related to this parameter. We clarify that the pebble flux proposed in this case is two orders of magnitude lower than the one proposed in a disk with , as the gas density is almost two orders of magnitude lower (in average during the disk lifetime), than the gas density in a disk with lower viscosity. We made this assumption following Lambrechts & Johansen (2014), as the pebble flux seems to scale with the gas density, and the gas density increases towards lower values of gas viscosity. However, we are interested in studying the impact of changing the pebble flux, thus the total pebble mass available for rocky planet assembly in future works to see the impact in the final planetary masses that could be achieved.
@@
We note that in all the scenarios it seems that the final configurations of the planetary systems are related mainly to the planet-disk interactions, and that tidal star-planet interactions are playing a second role in the dynamical evolution of the planets. In order to understand the importance of tidal interactions in the early rocky planet formation stages, we re-run the simulations from the scenarios that included pebble accretion in a gas disk with low viscosity, switching off the tidal effects. We have seen that the resulting planetary systems present the same range in mass and semi-major axis as the ones formed when including tidal effects. However, we found differences in the multiplicity of the systems and the frequency of planetary masses. When we neglected tidal interactions, we always formed planetary systems with higher multiplicity, thus, a bigger percentage of smaller planets per simulation. For a star of 0.1 M⊙, 70 of the simulations present planetary systems with more than 5 planets, while for the star of 0.3 M⊙ the percentage decrease to 60, and for a star of 0.5 M⊙ to 40. On the contrary, when including tidal effects for a star of 0.1 M⊙, 40 of the systems host more than 5 planets, for a star of 0.3 M⊙ just 10, and for a star of 0.5 M⊙ there is no planetary system with more than 5 planets. The higher multiplicity that we obtained when neglecting tidal effects is reflected in a higher percentage of smaller planets per scenario. We conclude that it is relevant to include tidal effects during rocky planet formation around M dwarfs, as it gives a more accurate estimate of the number of planets per system and the frequency of planetary masses. Interestingly, it seems that tidal effects enhanced the collisions among embryos, which is highly contributing to the final planetary masses.
@@ 6.2 Simulated planets vs observed exoplanets around M dwarfs
We highlight that in the formation scenarios in which we consider pebble accretion in a gas disk with low viscosity, the simulated planetary systems match the semi-major axis and masses of the observed low-mass close-in exoplanet sample less than 5 M⊕ for a star of 0.1 M⊙, between 2 and 10 M⊕ for a star of 0.3 M⊙ and between 3 and 17 M⊕ for a star of 0.5 M⊕.
@@
Moreover, we reproduce the inward shift in semi-major axis of close-in exoplanets around M dwarfs in agreement with Sabotta et al. (2021). As we can see from Figure 11, the planetary systems are located closer-in for the star of 0.1 M⊙, and further from the stars of 0.3 M⊙ and 0.5 M⊙. This is most probably related to the location of the inner edge of the disk for each stellar mass that we proposed (see Appendix A), which was based on the estimations of the stellar rotation period (Bouvier et al., 2014; Scholz et al., 2018). However, it seems we are overestimating innermost planets with periods less than 3 days. Thus, we will study the impact of the location of the inner gas disk radius in the location of the innermost planet population in future works.
@@
Furthermore, we found that most of the simulated systems are closely spaced. For stars close to 0.1 M⊙ this could be true if we take into account the observational biases towards close-in planets with masses less than 2 M⊕ (e.g. Ribas et al., 2023) and radius less than 0.9 R⊕ with transit technique (Ment & Charbonneau, 2023). However, it seems that for the more massive M dwarfs we are overestimating compact planetary systems, maybe due to a short integration time for the N-body simulations. Extending the simulations over a few Gyr would let us see if late collisions among planets could take place and if the star-planet tidal interactions have enough impact in the dynamical evolution of the systems to break such compact configurations. Unfortunately, this is out of the scope of this work due to the amount of CPU time required to run the
@@
N-body simulations. We note that other authors also find issues to break compact configurations in planetary systems around Sun-like stars (e.g. Izidoro et al., 2017, 2021). We point out that the mechanism that could make systems dynamically unstable is still under debate.
@@
In our simulations, we formed planets with masses between 0.35 and 10 M⊕ during the gas disk phase. Predominantly, rocky planets are expected to be smaller than  R⊕ and with masses less than 5 M⊕ to 10 M⊕, while bigger and more massive planets are more luckily to have substantial gaseous envelopes (e.g. Rogers, 2015). Moreover, planets with masses between 0.3 and 2 M⊕ are considered to have a pure rocky composition, while planets between 2 and 10 M⊕ are, most probably, water worlds Luque & Pallé (2022). We note that we could form planets with purely rocky cores just in the simulations that include an initial embryo distribution inside the snowline for a star of 0.1 M⊙. On the other hand, we could form water worlds planets (planets with around 50 of water in mass) around stars of 0.3 and 0.5 M⊙. In order to better constrain the final core composition of rocky planets it is necessary to include an initial sample of dry embryos located inside the snowline. Moreover, to better constrain the bulk composition of planets with masses higher than 5 M⊕, it will be important to take into account the fact that water worlds likely have steam atmospheres and this inflates the planetary radius with respect to most bulk models (Burn et al., 2024; Venturini et al., 2024).
@@
We note that there is a gap in planetary mass that we can see both in the confirmed exoplanet sample around M dwarfs and in the simulated planets from the scenarios that consider pebble accretion in a gas disk with low viscosity (see Figure 9). This gap seems to separate super-Earths and potential mini-Neptunes, which could be associated with the location of the radius valley for M dwarfs (Cloutier & Menou, 2020; Venturini et al., 2024). We note that in some of the simulations, we could form planets with masses bigger than the isolation mass during the gas disk phase, which could have retained a small fraction of gas in the envelope. In order to better constrain the core+envelope compositions of the simulated Super-Earths and potential mini-Neptunes, and confirm the relationship between the mass gap and radius valley, we need to include gas accretion in our model as well. This is something that we will study in future works and explore the probability that photoevaporation may not be necessary to explain the radius valley (e.g. Berger et al., 2023). In this case, it would also be interesting to study the effect of different disk lifetimes, as it will be related with the amount of gas available in the disk for planet gas accretion. However, this is something we did not test in this work as we did not consider planet gas accretion and most of the low-mass simulated planets formed achieve their final configurations early during the gas-disk phase.
@@
Finally, we would like to highlight the likely existence of small, inner dust traps due to the location of the outermost planets in our simulations that halts the inward pebble drift when reaching the isolation mass. Unfortunately, as the location of the outermost planets is around 0.1 au for a star of 0.1 M⊙ and 0.3 au for a star of 0.5 M⊙, such dust traps will remain unresolved, even with the James Webb Space Telescope (JWST). The MINDS JWST team could just detect dust traps that masses bigger than the isolation masses associated to each stellar mass generate indirectly by comparing H2O and CO2 fluxes as the CO2 snowline is further out (Grant et al., 2023). As our outermost planets are located inside both the H2O and CO2 snowline, JWST would not be able to confirm the existence of such small dust traps. However, such small-scale dust traps could exist as well, since they will still halt pebble drift and prevent mm-dust from disappearing entirely.
@@ 7 Conclusions
In this work we study the formation of low-mass close-in planets ( M⊕ and  au) around M dwarfs from compact dust disks ( au) with efficient radial drift to form planetary cores and with dust masses representative for initial disk dust masses from disk population models before radial drift effects () (Appelgren et al., 2023).
In order to do that, we developed different sets of N-body simulations that included a sample of embryos growing either by pebble accretion (see Section 3.4.1) or planetesimal accretion immersed in a gas disk for 10 Myr and exposed to planet-disk interactions (during the gas disk lifetime) as well as star-planet tidal interactions and general relativistic corrections that include the evolution of the stellar luminosity, radius and rotational period (see Appendix A and B). We proposed different formation scenarios, and in each of them we run 10 N-body simulations during 50 Myr. To start we considered a star of 0.1 M⊙ and assumed different initial distribution of embryos (embryos outside the snowline and embryos both inside and outside the snowline), two different viscosities of the gas disk ( and ) and two accretion mechanism (pebble accretion and planetesimal accretion). We also analyzed rocky planet formation around a star of 0.3 M⊙ and a star of 0.5 M⊙. In these cases we just changed the accretion mechanism, as we considered both samples of embryos that growth by pebble accretion and samples that growth by planetesimal accretion (see Section 3.1 for a detailed description of all the formation scenarios).
In Figure 4 we show a summary of the different formation scenarios in terms of the initial conditions for the disk and planetary cores together with the resulting simulated planetary masses in each case. We list below the main conclusions of this work:
@@
The final masses of the simulated close-in rocky planets around M dwarfs depend both on the pebble isolation mass and the collisions among planets.
The simulated planetary systems from the formation scenarios in which we consider efficient pebble accretion in a gas disk with low viscosity () are the ones that match the semi-major axis and masses of the observed low-mass close-in exoplanet sample around M dwarfs.
At least for a star of 0.1 M⊙, assuming a gas disk with high viscosity () and low pebble flux is not allowing the formation of either Earth-like or super-Earth like planets.
@@
To reproduce the observed rocky exoplanets with Earth-like composition, it is needed to start from an initial embryo population located inside the snowline that grows by an efficient pebble accretion.
From the good match that we found in our model between our simulated planets and the exoplanet sample of low-mass close-in exoplanets around M dwarfs in terms of masses, semi-major axis and multiplicity, we conclude that rocky planet formation around M dwarfs most probably took place in compact dust disks with efficient radial drift and in a gas disk with low viscosity. This work highlight the disk conditions needed for rocky planet formation to take place around very low mass stars.
@@ Acknowledgements.
- Banzatti et al. (2023) Banzatti, A., Pontoppidan, K. M., Carr, J., et al. 2023, arXiv e-prints, arXiv:2307.03846
@@
- Coleman et al. (2019) Coleman, G. A. L., Leleu, A., Alibert, Y., & Benz, W. 2019, arXiv e-prints, arXiv:1908.04166
@@
- Miguel et al. (2019) Miguel, Y., Cridland, A., Ormel, C. W., Fortney, J. J., & Ida, S. 2019, arXiv e-prints, arXiv:1909.12320
@@
- Venturini et al. (2024) Venturini, J., Ronco, M. P., Guilera, O. M., et al. 2024, arXiv e-prints, arXiv:2404.01967

### v3

@@ 6.1 Outcomes from the formation scenarios
The main conclusion from the different formation scenarios that we explored is the fact that close-in super-Earth formation is possible around M dwarfs in compact disks with low gas disk viscosity () when the core accretion is driven by an efficient pebble radial drift (pebble efficiency between 30 and 60). We found a linear relationship between the super-Earth masses and the stellar mass, in agreement with previous works (e.g., Liu et al., 2019). However, it is not just the isolation mass that sets the final planetary masses, but also the gravitational interactions among planets that lead to collisions among them, and thus an increment in their masses. Late collisions among planets become more relevant toward the more massive M dwarfs. As we can see in Section 4, only around a star of 0.5 M⊙ some simulated planets have masses higher than 10 M⊕. Interestingly, such planets had masses below 10 M⊕ while they were immersed in the gas disk. In every case, they reached higher masses after the gas dissipated from the disk, due to late collisions between planets.
@@
For the scenarios in which we study rocky planet formation around a star of 0.1 solar mass, it seems that the formation of Earth-like and super-Earth-like planets is independent of the initial location of the planetary embryos, in terms of their final masses and semi-major axis. However, the initial location of the embryos is important in terms of water content, as the scenario in which we considered a sample of embryos initially located inside the snowline is the only one that formed dry planets or planets with less than 10 of water in mass, more consistent with the observations (Luque & Pallé, 2022; Rogers et al., 2023). Even though we did not test the impact of different masses for the initial embryo distribution, we expect that it will not change significantly the final planetary mass range that we found: this is shown by Pan et al. (2024) when studying planet formation around a star of 0.1 M⊙ with , comparing the final planetary masses with both an initial equal mass embryos distribution and an embryo distribution with masses derivate from streaming instability (Liu et al., 2020). However, it would be interesting to see the dependence between the initial mass of the embryos with the final planetary masses in future works.
@@
On the other hand, assuming compact protoplanetary dust disks with higher viscosity () and lower pebble flux is not leading to an efficient planet formation, as all the simulated planets have masses . This is a crucial point as it seems that the migration speed of the embryos immerse in the gas disk as well as the pebble reservoir is directly related to this parameter. We clarify that the pebble flux proposed in this case is two orders of magnitude lower than the one proposed in a disk with , as the gas density is almost two orders of magnitude lower (in average during the disk lifetime), than the gas density in a disk with lower viscosity. We made this assumption following Lambrechts & Johansen (2014), as the pebble flux seems to scale with the gas density, and the gas density increases toward lower values of gas viscosity. However, we are interested in studying the impact of changing the pebble flux, and thus the total pebble mass available for rocky planet assembly in future works to see the impact in the final planetary masses that could be achieved.
@@
We note that in all the scenarios it seems that the final configurations of the planetary systems are related mainly to the planet-disk interactions, and that tidal star-planet interactions are playing a second role in the dynamical evolution of the planets. In order to understand the importance of tidal interactions in the early rocky planet formation stages, we re-run the simulations from the scenarios that included pebble accretion in a gas disk with low viscosity, switching off the tidal effects. We have seen that the resulting planetary systems present the same range in mass and semi-major axis as the ones formed when including tidal effects. However, we found differences in the multiplicity of the systems and the frequency of planetary masses. When we neglected tidal interactions, we always formed planetary systems with higher multiplicity, and thus a greater percentage of smaller planets per simulation. For a star of 0.1 M⊙, 70 of the simulations present planetary systems with more than 5 planets, while for the star of 0.3 M⊙ the percentage decrease to 60, and for a star of 0.5 M⊙ to 40. On the contrary, when including tidal effects for a star of 0.1 M⊙, 40 of the systems host more than 5 planets, for a star of 0.3 M⊙ just 10, and for a star of 0.5 M⊙ there is no planetary system with more than 5 planets. The higher multiplicity that we obtained when neglecting tidal effects is reflected in a higher percentage of smaller planets per scenario. We conclude that it is relevant to include tidal effects during rocky planet formation around M dwarfs, as it gives a more accurate estimate of the number of planets per system and the frequency of planetary masses. Interestingly, it seems that tidal effects enhanced the collisions among embryos, which is highly contributing to the final planetary masses.
@@ 6.2 Simulated planets vs observed exoplanets around M dwarfs
We highlight that in the formation scenarios in which we considered pebble accretion in a gas disk with low viscosity, the simulated planetary systems match the semi-major axis and masses of the observed low-mass close-in exoplanet sample less than 5 M⊕ for a star of 0.1 M⊙, between 2 and 10 M⊕ for a star of 0.3 M⊙ and between 3 and 17 M⊕ for a star of 0.5 M⊕.
@@
Moreover, we reproduce the inward shift in semi-major axis of close-in exoplanets around M dwarfs in agreement with Sabotta et al. (2021). As we can see from Figure 11, the planetary systems are located closer-in for the star of 0.1 M⊙, and further from the stars of 0.3 M⊙ and 0.5 M⊙. This is most probably related to the location of the inner edge of the disk for each stellar mass that we proposed (see Appendix A), which was based on the estimations of the stellar rotation period (Bouvier et al., 2014; Scholz et al., 2018). However, it seems we are overestimating innermost planets with periods less than three days. Thus, we will study the impact of the location of the inner gas disk radius in the location of the innermost planet population in future works.
@@
Furthermore, we found that most of the simulated systems are closely spaced. For stars close to 0.1 M⊙ this could be true if we take into account the observational biases toward close-in planets with masses less than 2 M⊕ (e.g., Ribas et al., 2023) and radius less than 0.9 R⊕ with transit technique (Ment & Charbonneau, 2023). However, it seems that for the more massive M dwarfs we are overestimating compact planetary systems, maybe due to a short integration time for the N-body simulations. Extending the simulations over a few Gyr would let us see if late collisions among planets could take place and if the star-planet tidal interactions have enough impact in the dynamical evolution of the systems to break such compact configurations. Unfortunately, this is out of the scope of this work due to the amount of CPU time required to run the
@@
N-body simulations. We note that other authors also find issues to break compact configurations in planetary systems around Sun-like stars (e.g., Izidoro et al., 2017, 2021). We point out that the mechanism that could make systems dynamically unstable is still under debate.
@@
In our simulations, we formed planets with masses between 0.35 and 10 M⊕ during the gas disk phase. Predominantly, rocky planets are expected to be smaller than R⊕ and with masses less than 5 M⊕ to 10 M⊕, while bigger and more massive planets are more luckily to have substantial gaseous envelopes (e.g., Rogers, 2015). Moreover, planets with masses between 0.3 and 2 M⊕ are considered to have a pure rocky composition, while planets between 2 and 10 M⊕ are, most probably, water worlds Luque & Pallé (2022). We note that we could form planets with purely rocky cores just in the simulations that include an initial embryo distribution inside the snowline for a star of 0.1 M⊙. On the other hand, we could form water worlds planets (planets with around 50 of water in mass) around stars of 0.3 and 0.5 M⊙. In order to better constrain the final core composition of rocky planets it is necessary to include an initial sample of dry embryos located inside the snowline. Moreover, to better constrain the bulk composition of planets with masses higher than 5 M⊕, it will be important to take into account the fact that water worlds likely have steam atmospheres and this inflates the planetary radius with respect to most bulk models (Venturini et al., 2020; Burn et al., 2024; Venturini et al., 2024).
@@
We note that there is a gap in planetary mass that we can see both in the confirmed exoplanet sample around M dwarfs and in the simulated planets from the scenarios that consider pebble accretion in a gas disk with low viscosity (see Figure 9). This gap seems to separate super-Earths and potential mini-Neptunes, which could be associated with the location of the radius valley for M dwarfs (Cloutier & Menou, 2020; Venturini et al., 2024). We note that in some of the simulations, we could form planets with masses bigger than the isolation mass during the gas disk phase, which could have retained a small fraction of gas in the envelope. In order to better constrain the core+envelope compositions of the simulated Super-Earths and potential mini-Neptunes, and confirm the relationship between the mass gap and radius valley, we need to include gas accretion in our model as well. This is something that we will study in future works and explore the probability that photoevaporation may not be necessary to explain the radius valley (e.g., Berger et al., 2023). In this case, it would also be interesting to study the effect of different disk lifetimes, as it will be related with the amount of gas available in the disk for planet gas accretion. However, this is something we did not test in this work as we did not consider planet gas accretion and most of the low-mass simulated planets formed achieve their final configurations early during the gas-disk phase.
@@
Finally, we would like to highlight the likely existence of small, inner dust traps due to the location of the outermost planets in our simulations that halts the inward pebble drift when reaching the isolation mass. Unfortunately, as the location of the outermost planets is around 0.1 au for a star of 0.1 M⊙ and 0.3 au for a star of 0.5 M⊙, such dust traps will remain unresolved, even with the James Webb Space Telescope (JWST). The MINDS JWST team could just detect dust traps that masses bigger than the isolation masses associated with each stellar mass generate indirectly by comparing H2O and CO2 fluxes as the CO2 snowline is further out (Grant et al., 2023). As our outermost planets are located inside both the H2O and CO2 snowline, JWST would not be able to confirm the existence of such small dust traps. However, such small-scale dust traps could exist as well, since they will still halt pebble drift and prevent millimeter-sized dust from disappearing entirely.
@@ 7 Conclusions
In this work we studied the formation of low-mass close-in planets ( M⊕ and au) around M dwarfs from compact dust disks ( au) with efficient radial drift to form planetary cores and with dust masses representative for initial disk dust masses from disk population models before radial drift effects () (Appelgren et al., 2023).
In order to do that, we developed different sets of N-body simulations that included a sample of embryos growing either by pebble accretion (see Section 3.4.1) or planetesimal accretion immersed in a gas disk for 10 Myr and exposed to planet-disk interactions (during the gas disk lifetime) as well as star-planet tidal interactions and general relativistic corrections that include the evolution of the stellar luminosity, radius, and rotational period (see Appendix A and B). We proposed different formation scenarios, and in each of them we ran ten N-body simulations lasting 50 Myr. To start we considered a star of 0.1 M⊙ and assumed different initial distributions of embryos (embryos outside the snowline, and embryos both inside and outside the snowline), two different viscosities of the gas disk ( and ), and two accretion mechanisms (pebble accretion and planetesimal accretion). We also analyzed rocky planet formation around a star of 0.3 M⊙ and a star of 0.5 M⊙. In these cases we just changed the accretion mechanism as we considered samples of embryos that grow by pebble accretion and samples that grow by planetesimal accretion (see Section 3.1 for a detailed description of all the formation scenarios).
In Figure 4 we showed a summary of the different formation scenarios in terms of the initial conditions for the disk and planetary cores together with the resulting simulated planetary masses in each case. We list below the main conclusions of this work:
@@
The final masses of the simulated close-in rocky planets around M dwarfs depend on the pebble isolation mass and on the collisions among planets.
The simulated planetary systems from the formation scenarios in which we considered efficient pebble accretion in a gas disk with low viscosity () are the ones that match the semi-major axis and masses of the observed low-mass close-in exoplanet sample around M dwarfs.
At least for a star of 0.1 M⊙, assuming a gas disk with high viscosity () and low pebble flux does not allow the formation of either Earth-like or super-Earth-like planets.
@@
To reproduce the observed rocky exoplanets with Earth-like composition, it is necessary to start from an initial embryo population located inside the snowline that grows by efficient pebble accretion.
From the good match that we found in our model between our simulated planets and the exoplanet sample of low-mass close-in exoplanets around M dwarfs in terms of masses, semi-major axis and multiplicity, we conclude that rocky planet formation around M dwarfs most probably took place in compact dust disks with efficient radial drift and in a gas disk with low viscosity. This work highlights the disk conditions needed for rocky planet formation to take place around very low-mass stars.
@@ Acknowledgements.
- Banzatti et al. (2023) Banzatti, A., Pontoppidan, K. M., Carr, J. S., et al. 2023, ApJ, 957, L22
@@
- Coleman et al. (2019) Coleman, G. A. L., Leleu, A., Alibert, Y., & Benz, W. 2019, A&A, 631, A7
@@
- Miguel et al. (2020) Miguel, Y., Cridland, A., Ormel, C. W., Fortney, J. J., & Ida, S. 2020, MNRAS, 491, 1998
@@
- Venturini et al. (2020) Venturini, J., Guilera, O. M., Haldemann, J., Ronco, M. P., & Mordasini, C. 2020, A&A, 643, L1
- Venturini et al. (2024) Venturini, J., Ronco, M. P., Guilera, O. M., et al. 2024, A&A, 686, L9